# Combining Teacher-Student with Representation Learning: A Concurrent Teacher-Student Reinforcement Learning Paradigm for Legged Locomotion

### v1

@@ III-A Reward Design
Feet distacne $r^{\text{fd}}$ $-100$
@@
In consideration of the higher degree of underactuation in bipeds compared to quadrupeds, maintaining base stability presents a more significant challenge. Accordingly, we have carefully decreased the penalty associated with Z-direction linear velocity. In parallel, we introduced a base orientation reward function that penalizes the roll and pitch angles of the base. This adjustment enables the bipedal robot to maintain a level posture across a variety of complex terrains as effectively as possible. Our observations indicate that the bipedal training strategy tends to favor foot placements closer to the center-line to minimize torque on the center of mass. However, this behavior increases the risk of leg collisions. To mitigate this issue, we have incorporated a feet distance penalty function, denoted as $r^{\text{fd}}$:
@@
Due to the point-foot bipedal robot’s lack of sole structures, it is unable to maintain static standing balance. Therefore, we referred to the method in [26] and added periodic rewards to encourage the generation of a regular trotting gait for self-balancing. The rewards $r^{\text{ff}}$ and $r^{\text{fv}}$ penalize the foot contact forces during the stance phase and foot velocities during the swing phase, respectively, allowing the agent to learn specified contact patterns.
@@
III-B Linear Velocity Estimator
We consider the linear velocity of the base is part of privileged state $\boldsymbol{s}_{t}$ instead of proprioceptive observations $\boldsymbol{o}_{t}$, due to the challenges associated with obtaining accurate and reliable velocity estimation by using proprioceptions in real-world. Quadrupedal robots have enough legs in contact with the terrain that explicit estimates of velocity are not indispensable. However, point-foot bipedal robot lacks enough contact information. This can lead to a decrease in the ability of bipedal robots to navigate complex terrain. Therefore, we adopt the method proposed in  [12] and additionally integrate a velocity estimator for bipedal robots, which uses observation sequence $\boldsymbol{o}^{\text{s}}_{t-H:t}$ to estimate base linear velocity and provide it to the policy network.
III-C Environment Setup
We use IsaacGym simulator [28] to train 8192 parallel agents on different terrains, each group contains 4096 agents. The training process needs 3000 iterations to get capability of traversing challenging terrains, which takes about 1 hour 45 minutes of wall clock time on NVIDIA RTX 4090. But its performance will continue to improve over an extended period of training. Each episode lasts no longer than 20s, corresponding to 1,000 time steps with a control frequency of 50 Hz. Episodes were terminated upon reaching the maximum time or robot fall-over. Joint PD controller parameters are $k_{\text{p}}=20.0$, $k_{\text{d}}=0.5$ for A1, $k_{\text{p}}=40.0$, $k_{\text{d}}=1.0$ for Aliengo, $k_{\text{p}}=40.0$, $k_{\text{d}}=1.5$ for biped. The algorithm performed an iteration every 24 time steps. Hyper-parameters for training are presented in Table III, and adaptation of the learning rate is similar to [13].
@@ III-C Environment Setup
Batch size $4096\times 24$
Mini-batch size $4096\times 6$
@@
To achieve robust locomotion in various terrain, it is crucial to implement a proper training curriculum strategy. We adopt a terrain curriculum similar to [13]. At the beginning of training, all robots are assigned to four terrain types with the lowest difficulty, including slopes, rough slopes, stairs, and discrete obstacles shown in Fig. LABEL:fig:sim_terrain. The robots are moved to more difficult terrains once they successfully mastering velocity tracking.
@@
To compensate for the gap between simulation and the real-world, we randomize mass of base and legs, CoM of base, the friction and restitution between the rigid body and the ground, the PD gains and motor strength and position offsets of joints, and action delay. The details of the randomization are presented in Table IV.
@@ IV Results
IV-A Latent Space Analysis
To elucidate the acquired representations, we analyzed the structure of the latent space. We used the proprioceptive encoder to obtain latent representations on different terrains (sopes, rough slopes, stairs, and discrete obstacle) in simulation, and compare the representations encoded by different robots. We conducted t-distributed stochastic neighbor embedding (t-SNE) on the latent representations for the above terrains, as shown in Fig. 6. We found that the latent representations were distributed distinctly for different terrains and thus carried sufficient information about the features of terrains. Moreover, the t-SNE plot demonstrates that quadruped robots possess a stronger capability to perceive terrain features compared to bipedal robots, as they have more legs to make contacts with the environment.
Figure 6: The t-SNE visualization of Latent Representations from Encoder for A1 and Biped
IV-B Evaluation Results
@@
  • Oracle: Policy full state was trained by PPO with.
@@
  • EstimatorNet: Policy was concurrently trained with an estimator network estimates body velocity and feet height similar to  [12].
@@
  • Two-stages teacher-student: The proposed method trained in two stages where student policy was trained by supervised learning as original teacher-student learning framework.
@@
For a fair comparison, all the methods above were trained under the same training configuration detailed in Section III-C, and use the same network scale and random seeds. Our evaluations will employ the policy at the 5000th iteration. Specifically, for the two-stages teacher-student method, the policy is obtained by 3000 iterations for teacher and 2000 iterations for student.
We compared improvements in terrain level and tracking reward during the training process. The results are shown in Fig. 7, in which the curves are averaged over 5 seeds. The shaded area represents the standard deviation across seeds. The Baseline and EstimatorNet models exhibit comparable capabilities in traversing varied terrains and similar training velocities. However, the Baseline shows significant variability with different random seeds. Oracle model with access to privileged information achieves the highest training velocity. In the later stages of training, to minimize collisions and other penalties for higher discounted returns, it slightly compromises command tracking under specific conditions, as indicated by a minor decline in the terrain level on the graph. The training velocity of our teacher model is marginally slower than that of the Oracle, and the student model lags slightly behind the teacher. Nevertheless, both eventually attain equivalent capabilities.
Figure 7: Learning curves of average terrain level.
We evaluated the velocity tracking performance in simulation under various terrain conditions by distributing 8192 robots evenly among four types of terrains. Linear velocity commands were uniformly sampled from [-1.0, 1.0] m/s. Tracking errors are quantified using the norms $\|\boldsymbol{v}_{xy}^{\text{cmd}}-\boldsymbol{v}_{xy}\|_{2}$. To minimize the impact of randomness during the training process, the results presented in the Table V are the average outcomes of policies trained with five different seeds. It shows that the tracking capability of our method outperforms the other methods on all terrains.
TABLE V: Velocity Tracking Error
Method Slope Rough Stairs Obstacles
Baseline 0.119 0.165 0.195 0.132
EstimatorNet 0.111 0.158 0.193 0.125
Two-stages teacher-student 0.103 0.141 0.138 0.113
Ours 0.098 0.128 0.133 0.105
To evaluate the robustness of policy, we perturbed robots in the simulation under various terrains by applying random force to the robot’s body and record their survival rates. Specifically, the direction of the push was chosen randomly, with the applied force inducing a velocity change of approximately 2m/s of the robots. Results in Table VI show that our method is more robust in challenging terrains.
TABLE VI: Push Survival Rate in %
Method Slope Rough Stairs Obstacles
Baseline 97.13 93.78 94.72 95.29
EstimatorNet 95.83 93.21 94.05 95.01
Two-stages teacher-student 96.86 94.04 94.87 95.92
Ours 96.34 94.50 95.24 95.58
Figure 8: Aliengo steps over a moving platform
@@
Figure 9: A1 locomotion with stairs and missing step adaptation
@@
Figure 10: Biped locomotion with stairs
IV-C Real-World Experiments
@@
As the platform commences sliding upon the robot’s ascent, the robot reacts to this motion. Its hind legs promptly engage, adjusting to maintain balance in response to the new dynamics introduced by the moving platform. This demonstrates the robot’s reactive balance capabilities and its proficiency in stabilizing itself amid changing environmental conditions. As the robot descends from the platform, the placement of its left hind leg is dynamically adjusted in relation to the robot’s overall posture and the positioning of its other legs. During its swing phase, it selects a more forward foothold to align with the robot’s overarching motion, resulting in an S-shaped swing trajectory.
Fig. 9 provided visual and data analysis illustrates quadruped robot’s motion as it encounters stairs and missing step. When the robot encounters a step while walking, its right front leg is quickly lifted after being stumbled, allowing it to step onto the stair. The robot then smoothly reaches the top and proceeds toward the edge until its front leg steps into empty space, initiating a fall. Immediately sensing the change in terrain, the robot reacts swiftly, extending its front legs to seek additional support and keeping its center of mass within a safe zone determined by the stance of its limbs. Once the front legs make contact, the hind legs follow, enabling the robot to safely return to the ground.
Fig. 10 demonstrates the effectiveness of our policy applied to a more challenging bipedal platform. The policy enables the robot to maintain stable forward locomotion on flat surfaces. When the right leg strikes the edge of the stair, the velocity of the knee joint rapidly drops to zero to prevent loss of balance due to inertia, and lifts upwards to navigate over the obstacle of the stair. Then, when the left leg ascends the stair, the robot briefly adopts a double-support stance to secure its base stability. After perceiving the height of the initial step, the biped transitions to the next level smoothly, avoiding collisions and demonstrating effective stair-climbing.
V Conclusion

### v2

@@ III-A Reward Design
Feet distance $r^{\text{fd}}$ $-100$
@@
In consideration of the higher degree of underactuation in bipeds compared to quadrupeds, maintaining base stability presents a more significant challenge. Accordingly, we have carefully decreased the penalty associated with the linear velocity in the Z direction and introduced a base orientation reward to penalize the base roll and pitch angles. This adjustment enables the bipedal robot to maintain a level posture across a variety of complex terrains as effectively as possible. Our observations indicate that the biped tends to favor foot placements closer to the center-line to minimize torque on the center of mass. However, this behavior increases the risk of leg collisions. To mitigate this issue, we have incorporated a feet distance penalty function, denoted as $r^{\text{fd}}$:
@@
Due to the lack of sole structures, the point-foot bipedal robot is extremely difficult to maintain static standing balance. Therefore, we refer to the method in [27] and add periodic rewards to encourage the generation of a regular periodical gait for self-balancing. The rewards $r^{\text{ff}}$ and $r^{\text{fv}}$ penalize the foot contact forces during the stance phase and foot velocities during the swing phase, respectively, allowing the agent to learn specified contact patterns.
@@
III-B Environment Setup
We use IsaacGym simulator [29] to train 8192 parallel agents on different terrains, To keep the leadership role of the teacher in skill learning, we distribute the agents among the two groups with a ratio of 3:1, i.e., teacher group and student group consist of 6144 and 2048 agents, respectively. A better expression: It takes around 3000 iterations for the policy to acquire the ability of handling challenging terrains like stairs, which corresponds to around 105 minutes of wall clock time using a Nvidia RTX 4090 for training The policy’s performance will continue improving with further training. During the training process, we set the maximum time duration for each episode to be 20 seconds, corresponding to 1,000 time steps with a control frequency of 50 Hz. All episodes were terminated upon reaching the maximum time duration or upon experiencing a robot fall-over. The joint PD controller parameters are set to be $k_{\text{p}}=20.0$, $k_{\text{d}}=0.5$ for A1, $k_{\text{p}}=40.0$, $k_{\text{d}}=1.0$ for Aliengo, $k_{\text{p}}=40.0$, $k_{\text{d}}=2.5$ for P1. The length of the observation sequence $\boldsymbol{o}^{\text{s}}_{t-H:t}$ is set to 5 for both quadruped and bipedal policies, following the settings reported in [25] The algorithm performed an iteration every 24 time steps. Hyper-parameters for training are presented in Table III, and adaptation of the learning rate is similar to [15].
@@ III-B Environment Setup
Batch size $2048\times 24$
Mini-batch size $2048\times 6$
@@
To achieve robust locomotion in various terrain, it is crucial to implement a proper training curriculum strategy. We adopt a terrain curriculum similar to [15] and have selected four terrain types for our training: slopes, rough slopes, stairs, and discrete obstacles, as shown in Fig. 3. The slopes and rough slopes have gradients ranging from $0^{\circ}$ to $26.57^{\circ}$, with rough slopes additionally containing uniform noise ranging from 5 cm to 17 cm. The stairs have heights ranging from 5 cm to 23 cm, and the discrete obstacles have heights ranging from 5 cm to 24 cm. Each terrain type is divided into difficulty levels from 0 to 9, evenly distributed within the specified difficulty ranges. At the beginning of training, all robots are assigned the lowest difficulty level of these four types of terrain. The robots are moved to more difficult terrains once they have traversed the current area.
(a) slopes
(b) rough slopes
(c) stairs
(d) discrete obstacles
Figure 3: Terrains in simulation.
@@
To compensate for the gap between simulation and the real world, we randomize inertial parameters such as the mass of the base and legs, the CoM of the base, the friction and restitution between the rigid body and the ground, the PD gains and motor strength, and action delay. The details of the randomization are presented in Table IV.
@@
Figure 4: Learning curves of average terrain level.
@@ IV Results
IV-A Evaluation Results
@@
  • Oracle: Policy receives the encoded privileged state as input and was trained by PPO.
@@
  • EstimatorNet: Policy was concurrently trained with an explicit estimator network estimating body velocity and feet height similar to [14].
@@
  • Two-stage teacher-student (T-S): The proposed method trained in two stages where student policy was trained by supervised learning using latent reconstruction loss and action imitation loss, following the original teacher-student learning framework as described in [16].
  • Regularized Online Adaptation (ROA): A single-stage training method presented in [24].
@@
For a fair comparison, all the methods above were trained within asymmetric actor-critic framework under the same training configuration detailed in Section III-B, and use the same network scale and random seeds. Our evaluations will employ the policy at the 5000th iteration. Specifically, for the two-stage teacher-student method, the policy is obtained by 3000 iterations for teacher and 2000 iterations for student. For ROA, we obtain the same switch period and regularization curriculum as described in [24].
Figure 5: Evaluation of average tracking error in four types of terrains. Linear velocity commands were uniformly sampled from $[-1.0,1.0]$ m/s.
Figure 6: Evaluation of push recover in four types of terrains. Push by applying force inducing a velocity change of approximately 2.5m/s of the robots.
We compared improvements in terrain level during the training process. The results are shown in Fig. 4, in which the curves are averaged over 5 seeds. The shaded area represents the standard deviation across seeds. The terrain level curve records the average terrain level of all agents at each moment during the training process. It serves as a statistical measure of the agents’ overall mobility capabilities. Fig. 4 shows that teacher in CTS exhibits performance nearly identical to the Oracle, indicating that synchronous training with the student does not impair the teacher’s performance. Student in CTS performs slightly worse than the teacher but still outperforms the student policy trained in two-stage teacher-student by imitating teacher. The baseline, which does not directly utilize privileged information in agent decision-making, exhibits similar final performance to EstimateNet. However, due to the guidance provided by supervised learning signals for body velocity and feet height, EstimateNet shows faster initial learning compared to the baseline. Due to the policy-switching characteristic of ROA during training, the terrain level curve does not effectively reflect the mobility capabilities of individual policies and lacks reference value; thus, it has been excluded from consideration.
Figure 7: Aliengo steps over a moving platform.
We evaluated velocity tracking performance under various terrain conditions by distributing 8192 robots evenly among four types of terrains. Velocity commands were uniformly sampled from $[-1.0,1.0]$ m/s. Tracking errors are quantified using the metric $\|\boldsymbol{v}_{xy}^{\text{cmd}}-\boldsymbol{v}_{xy}\|_{2}$. Fig. 5 presents the average tracking error (y-axis) of different training methods under four types of terrains. The points represent the average values of the policies trained under 5 different seeds, while the lines indicate the standard deviation. It shows that student of CTS achieves a reduction in velocity tracking error compared to the student trained in two-stage teacher-student, with improvements of 17.85% on slopes, 19.12% on rough slopes, 7.9% on stairs, and 21.85% on discrete obstacles. On the stair terrain, the velocity tracking performance of the baseline and EstimateNet is relatively poor compared to other methods. We believe this is because the stair terrain is the most challenging for maintaining good velocity tracking among all terrains. Therefore, it requires the terrain information from the privileged information more significantly, which explains why the three methods directly guided by privileged information exhibit better velocity tracking performance on the stair terrain.
@@
Figure 8: Quadruped locomotion with stairs and missing step adaptation. The curves record the joint angles of the robot’s right front leg.
@@
Figure 9: Biped locomotion with stairs. The curves record the joint angles of the robot’s right leg.
To evaluate the robustness of policies, we apply random force to the robot’s body and record their survival rates on various terrains. Specifically, the direction of the push was chosen randomly, with the applied force inducing a velocity change of approximately 2.5m/s of the robots. Results in Fig. 6 shows CTS demonstrates greater robustness across various terrains compared to the two-stage teacher-student. The survival rates under random push for the four types of terrain are higher by 5.04%, 6.47%, 4.76%, and 4.57%, respectively. The staged training student exhibits the poorest robustness against random pushes. This may be due to its sole training objective of imitating the teacher, without considering the reinforcement learning goal of maximizing expected discounted return. Due to the information gap, the student cannot perfectly mimic the teacher, ultimately leading to a lack of emphasis on maintaining balance and making it more prone to being toppled by pushes. The figure also shows that EstimateNet performs quite well across various terrains, which may benefit from its explicit estimation of the base linear velocity. We believe that adding a similar mechanism to our method will help further improve robustness.
Based on the results of the comparative evaluation, we believe that in training policies that can only access proprioceptive observations, directly incorporating privileged information to aid decision-making, along with reinforcement learning aimed at maximizing the expected discounted return, helps in achieving robust and high-performance policies.
IV-B Real-World Experiments
@@
As the platform commences sliding upon the robot’s ascent, the robot’s hind legs promptly engage, adjusting to maintain balance in response to the new dynamics introduced by the moving platform. This demonstrates the robot’s reactive balance capabilities and its proficiency in stabilizing itself amid changing environmental conditions. As the robot descends from the platform, the placement of its left hind leg is dynamically adjusted in relation to the robot’s overall posture and the positioning of its other legs. During its swing phase, it selects a more forward foothold to align with the robot’s overarching motion, resulting in an S-shaped swing trajectory.
Fig. 8 provides visual and data analysis illustrating the quadruped robot’s motion as it encounters stairs and missing steps. The curves in the figure represent the angular trajectories of the hip and knee joints of the robot’s right front leg. When the robot encounters a step while walking, its right front leg is quickly lifted after being stumbled, allowing it to step onto the stair. The robot then smoothly reaches the top and proceeds toward the edge until its front leg steps into empty space, initiating a fall. Immediately sensing the change in terrain, the robot reacts swiftly, extending its front legs to seek additional support and keeping its center of mass within a safe zone determined by the stance of its limbs. Once the front legs make contact, the hind legs follow, enabling the robot to safely return to the ground.
Fig. 9 demonstrates the effectiveness of our policy applied to a more challenging bipedal platform. The curves in the figure represent the angular velocity trajectories of the hip and knee joints of the robot’s right leg. The policy enables the robot to maintain stable forward locomotion on flat surfaces. When the right leg hits the edge of a step during the swing phase (as indicated by the instantaneous decrease in the angular velocity of the hip joint in the figure), the swing trajectory is forced to change, causing the foot to land prematurely. At this point, the policy has already detected the presence of the obstacle through proprioceptive observation. Subsequently, the left leg is raised directly to overcome the obstruction. After the left leg lands, the policy perceives the approximate height of the stair, allowing subsequent swing trajectories to have sufficient height to smoothly ascend the stair.
V Conclusions and Future Works